OpenStudy (anonymous):

can some one help me solve y-2x-6=0 and y/2-3=x solving for 2 variables

OpenStudy (anonymous):

y-2x-6=0 y/2-3=x put x=y/2 -3 into y-2x-6=0 y-2(y/2 -3)-6=0 y-y+6-6=0 0=0 Therefore, y can be any real numbers x is the same Ans: all real solution